Maddie has decided the take a job as the new mascot of the MN United Loons. For her new stunt, she will stand on the lower deck and kick a ball, straight up, to land in a bucket held by an assistant in the edge of the upper deck. If she kicks the ball straight up with a velocity of 24 m/s and it lands in the bucket after 2.5 seconds, how high is the upper deck?


1
Expert's answer
2020-12-21T11:00:03-0500

"\text{We have the movement of the body thrown vertically up}"

"\text{Formula for uniformly accelerated motion,we take that }a= -g"

"h = v_ot-\frac{gt^2}{2}=24*2.5-\frac{9.8*2.5^2}{2}=29.375"

Answer:29.375 m